# CirrMRI600+: Large Scale MRI Collection and Segmentation of Cirrhotic Liver

### v1

@@ 3 Proposed Dataset: CirrMRI600+
4 Experiments
4.1 Experimental settings
Baseline methods. We carefully selected several baseline methods from CNN-family (e.g., VNet [37], AttentionU-Net [22]) and from transformer-based methods (for example, nnUNet [31], Swin UNetR [32], nnFormer3D [36], LinTransUNet [35] and TransUNet3D [38].
Implementation details. We trained our models using PyTorch 2.2.2 with CUDA 11.2. The learning rate was initially set to 0.0001 and was gradually decreased using the Cosine Annealing Scheduler. We use the BCE-Dice loss with the AdamW optimizer with a batch size of 4. The models were trained for 500 epochs with an early stopping patience of 50 epochs to prevent overfitting. The learning rate decay was set at 0.001 after every 10 epochs. To accelerate training, we leveraged two Nvidia A6000 GPUs, each with 48GB of memory, and utilized PyTorch’s Distributed Data Parallel to distribute a batch of 4 to each GPU. We resized every volume to uniform spatial dimensions of for generalizability.
4.2 Benchmark Results
@@
Quantitative analysis on CirrMRI600+T1W: Table 2 presents a comprehensive evaluation of 11 SOTA 3D segmentation networks on the CirrMRI600+T1W dataset. As summarized, nnSynergyNet3D achieved the highest overall performance with mIoU of 84.51, DSC of 87.89%, HD95 of 21.04 mm, and precision of 88.72%. nnSynergyNet3D performed better because of its synergistic and auto-configured continuous and discrete representation, allowing the model to capture fine and coarse features along with long-range dependencies due to its Transformer-inspired design. LinTransUnet and nnFormer3D demonstrated comparable performance in capturing cirrhotic liver tissue and its boundaries. Their performances were attributed to their Transformer based design with auto configuration, which enabled the models to learn and adapt to the liver’s varying shape and complex boundaries. Again, this highlighted the importance of long-range dependencies. Conversely, nnUNet3D demonstrated slightly poorer performance, underscoring the significance of Transformer-based representations for cirrhotic liver segmentation. Models like SwinUNeTr, TransBTS, and TransUNet3D do not significantly surpass CNN-based models such as nnUNet and SynergyVNet3D, showing the importance of auto-configuration and hybrid CNN-Transformer-based models.
Quantitative analysis on CirrMRI600+T2W: Table 3 presents evaluation of SOTA 3D segmentation networks on the CirrMRI600+T2W dataset. We observed similar results to those in the T1W segmentation results. nnSynergyNet3D has a superior DSC value of 86.51%, the lowest HD of 24.19 mm, and the lowest ASDD value of 3.96 mm. nnFormer3D and nnUNet3D are the other two competitive networks. The models such as SwinUNeTr, TransBTS, and TransUNet3D do not significantly surpass CNN-based models such as nnUNet and SynergyVNet3D, emphasizing the importance of auto-configured and hybrid CNN-Transformer-based models for achieving competitive segmentation results.
@@
Qualitative analysis on CirrMRI600+T1W and CirrMRI600+T2W: Fig. 3 shows qualitative results for T1W and T2W samples, respectively. These results demonstrated that existing models perform well in segmenting cirrhotic liver under mild conditions. As highlighted by white boundaries, these models suffer under moderate-to-severe cases due to the poor texture of MRIs caused by cirrhosis scarring. nnSynergyNet3D consistently performed well even for advanced cirrhotic livers compared to other SOTAs such as nnUNet, TransUNet, and Attention-UNet. This is likely because of its auto-configured, hybrid, and synergistic nature, enabling it to capture the texture of the cirrhotic liver more effectively.
@@
Cross-domain analysis: Table 4 shows results of segmentation networks pretrained on CirrMRI600T1W MRIs and tested on the part with CT scans pertaining to LiTS dataset. Results indicated that training pre-trained models with auto-configured architectures such as nnSynergyNet3D, on the CirrMRI600T1W dataset improves generalization to liver CTs. This further provides valuable insight and promise towards building auto-configured synergistic models to be pre-trained on our proposed dataset. Such an approach could potentially reduce the workload of clinicians working on both CT and MRI modalities.
Method mIoU Dice HD95 Precision Recall ASSD nnUNet3D [[31]] 80.02 81.19 29.78 80.98 83.21 6.52 nnFormer3D [[36]] 80.41 82.19 29.19 82.76 85.33 5.91 TransUNet3D [[38]] 76.87 77.89 33.78 78.87 80.02 7.08 SwinUNITR [[32]] 78.89 79.12 30.98 77.98 80.03 6.67 TransBTS [[40]] 79.11 81.11 29.48 79.91 81.19 6.01 nnSynergyNet3D [[43]] 81.88 83.38 29.01 81.15 86.02 5.55
5 Discussion
Liver cirrhosis is a global health issue [46]. While there are existing datasets for healthy liver subjects ( e.g., CHAOS [38]) and MRI segmentation datasets (e.g., Duke Liver segmentation dataset [42]), none of them are specifically designed for liver cirrhosis. The anatomy of a healthy liver is different compared to cirrhosis. There are complex challenges in cirrhotic liver cases, including texture distortion, shape variation, and boundary ambiguity. These factors pose significant obstacles to the development of robust deep learning methods. With CirrMRI600+, we have already achieved strong performance for T1W (DSC of 87.89%) and T2W (86.51%) with nnSynergyNet3D [43]. The team of four clinicians verified all the volumetric scans, but only one radiologist double-checked them. Still, we obtained high inter-observer agreement (kappa scores of 0.89 for T1W and 0.87 for T2W). We believe that using novel architectural designs and data augmentation techniques specifically for cirrhotic liver cases can further improve performance and generalizability.
While CirrMRI600+ offers a significant leap forward in liver cirrhosis segmentation research, it’s important to acknowledge its limitations. One such limitation is its single-center nature. Ideally, a truly comprehensive dataset would encompass data from multiple medical institutions, capturing the potential variations in acquisition protocols and patient populations encountered in real-world clinical settings. However, CirrMRI600+ addresses this limitation by prioritizing other key strengths. The dataset was very carefully curated, ensuring the highest quality annotations and ground truth segmentation for all included scans. Furthermore, CirrMRI600+ boasts a rich heterogeneity in disease states within the cirrhotic population. By carefully selecting a diverse range of cirrhotic liver presentations, the dataset ensures generalizability to a broader spectrum of cirrhosis cases.
Another weakness was that more than 95% of the T1W images were contrast enhanced and not all possible contrast types and phases of contrast administration were represented in the dataset. Although the cross-modality testing shows strong generalizability, small annotations might have been missed, and our annotations might not be correct pixel precise for all the cases. Other modalities, such as computed tomography (CT) scan, ultrasound, elastography, and magnetic resonance elastography, were absent besides MRI.
6 CirrMRI600+: Additional Information
Motivation for experiments using 2D slices for CirrMRI600+ T2W:
The T2W images were predominantly acquired using accelerated techniques such as half-fourier single shot turbo-spin-echo (HASTE). The T2W images were acquired in a slice-by-slice fashion in the axial plane (the physics of HASTE generates pseudo-3D volumes). Due to the long echo times required for T2W image acquisition, motion artifacts in the craniocaudal plane from diaphragm movement during respiration are expected. Due to this, we trained T2W-based segmentation models using both volumetric data and 2D planar data. Additionally, we show the comparisons between the different models. By including the experiments on the 2D dataset, we want to develop a robust algorithm that could work well with motion artifacts that might have affected the 3D volumetric scans. We believe that for segmentation models using T2W images, analysis and segmentation using 2D images is an appropriate choice given the expected motion artifacts arising from the acquisition techniques.
7 Results for CirrMRI600+T2w (slice-by-slice)
In this section, we present implementation details, baselines used for comparisons, and results.
@@ 7 Results for CirrMRI600+T2w (slice-by-slice)
7.1 Implementation details
We performed all our 2D experiments on the NVIDIA RTX A6000 GPU system for computing. The network is configured to train with a batch size 16 and a learning rate set to 1e-4. We train all the models for 500 epochs to fine-tune the network parameters adequately with an early stopping patience of 50. To enhance the performance of different networks, we used a combined loss function that consists of binary cross-entropy and dice loss. We use Adam optimizer for all the experiments for efficient parameter updates. We use the same evaluation metrics such as mIoU, dice coefficient, HD95, precision, recall and ASDD for result comparison.
7.2 Baseline methods
Here, we provide six baselines for benchmarking on CirrMRI600+T2w 2D datasets. For the experiments, we chose architectures such as UNet, AttentionUNet, nnUNet2D, TransUNet, SynergyNet, and MedSegDiff. Unless otherwise specified, we followed the default configurations provided in their released codebases. By including a diverse range of algorithms, from encoder-decoder-based networks such as UNet and AttentionUNet to more specialized and advanced architecture such as nnUNet2D (that uses deep supervision), TransUNet (transformer based), SynergyNet, and MedSegDiff (diffusion based model), that covers different approaches and techniques in medical image segmentation.
7.3 Quantitative and qualitative results
Table 5 presents the quantitative results comparison of different methods across various metrics on CirrMRI600+T2w 2D datasets. Among the UNet-based methods, SynergyNet emerges as the top performer across all metrics for the liver cirrhosis dataset. SynergyNet obtains a mIoU value of 0.7383, dice coefficient of 0.7592, HD95 30.94, precision of 0.7882, recall of 0.8222 and ASDD of 7.55. This suggests the superior capability of SynergyNet in accurately segmenting liver cirrhosis tissues from T2-weighted MRI scans. Interestingly, MedSegDiff, a method based on diffusion models, outperforms all other approaches. MedSegDiff obtains an mIoU of 0.7489, dice coefficient of 0.7667, HD95 of 30.89, precision of 0.7789, recall of 0.8192 and ASDD of 7.34. MedSegDiff has better mIoU, dice, HD, and ASDD values, whereas SynergyNet has better precision and recall values.
MedSegDiff’s superior performance can be attributed to its utilization of a diffusion process combined with a Fourier-based noise predictor. This unique combination enables the model to capture finer details and subtle variations in the input data more effectively, resulting in improved segmentation accuracy. These results underscore the potential of diffusion models in T2W MRI liver cirrhosis segmentation. We hope that our baseline experiments can encourage other researchers in the field, which can lead to developing novel methods for liver cirrhosis segmentation tasks.
Figure 4 shows the qualitative results of different methods on CirrMRI600+T2w datasets. From the qualitative results, it can be observed that models such as UNet produce under-segmentation, whereas nnUNet produces over-segmentation. Synergynet also produces shows under and over-segmentation for diverse cases. However, MedSegDiff is better at handling complexity.
7.4 Dataset availability and benchmarking algorithm
Our dataset is publically available at https://osf.io/cuk24/. We have subdivided the dataset into three folders, namely, CirrMRI600+T1w, CirrMRI600+T2w and CirrMRI600+T2w 2D. Researchers can download our dataset with a single click. Our dataset has been made publicly available under a Creative Commons Attribution-ShareAlike 4.0 International (CC-BY-SA 4.0) license, which allows for data sharing, redistribution, and academic and commercial usage (with written permission required for commercial purposes). We have provided the recommended train, valid and test split for the dataset. We also provide the information about the baseline algorithms here 222https://github.com/NUBagciLab/CirrMRI600Plus so that other researchers can reproduce our methods.
7.5 Author statement
We, the authors, hereby confirm that we have thoroughly verified the origins and licensing of the CirrMRI600+ dataset. We will bear full responsibility for any potential rights violations, including but not limited to intellectual property rights, privacy rights, or any other legal or ethical concerns arising from releasing and using the CirrMRI600+ dataset. We also confirm that all the necessary measures have been taken to ensure data anonymization and de-identification of sensitive information in the dataset (metadata).
We ensure that our dataset will be available in the long term. Upon acceptance of the work, we will also host our dataset on a platform such as https://zenodo.org/, where it offers to have a separate DOI (for example, 10.xxxx/zenodo.xxxxxx).
7.6 Raw data and Meta data information
The CirrMRI600+ dataset comprises volumetric scans of T1-weighted MRI, T2-weighted MRI, and 2D scans for T2-weighted MRI for cirrhotic livers. We have provided corresponding labels for each volumetric scan and 2d scans. We do not provide details such as image numbers or raw data information from the original dataset as they are not relevant for public release. The provided volumetric scans and their corresponding segmentation masks are sufficient for researchers to develop novel methodologies and compare the existing deep learning algorithms.
Additionally, we do not release information such as patient name, patient ID, date of birth, other patient ID, age, weight, accession number, referring physician name, device serial number, accession number, exact study date of each patient, series date, acquisitions date, etc. to preserve patient anonymity. This information would ensure data safety.
8 Conclusion
@@
We presented CirrMRI600+, a novel dataset specifically designed for cirrhotic liver segmentation. This dataset addresses a critical need in the field of deep learning for liver disease analysis. It comprises a collection of T1-weighted (T1W) and T2-weighted (T2W) volumetric MRI scans from 339 patients diagnosed with cirrhotic liver disease, along with corresponding ground truth segmentation masks for each scan. To ensure high-quality annotations while maintaining efficiency, CirrMRI600+ leverages a semi-automated approach in ground trusth construction. Deep learning model predictions served as a starting point, followed by refinement from radiologists. This approach enabled radiologists to focus on potential algorithm failures and refine the output masks more efficiently, saving valuable annotation time and cost. Rigorous benchmarking using the nnSynergyNet3D method on both T1W and T2W scans demonstrates promising performance, achieving Dice Similarity Coefficients (DSC) of 87.89% and 86.51%, respectively. CirrMRI600+ lays the foundation for a future multi-organ dataset. Our vision is to expand the annotations to include additional organs such as kidneys, spleen, pancreas, and major vessels, creating a comprehensive resource for abdominal organ segmentation tasks for multimodal MRIs.

### v2

@@ Abstract
Data Record
All data records collected in this study can be found in the OSF servers [25] (DOI:10.17605/OSF.IO/CUK24). All the medical images (MRI T1W and T2W) are stored using digital imaging techniques in the NIfTI format. The segmented images of the liver and the matched original medical images were stored in the NIfTI format after segmentation. The CirrMRI600+ dataset consists of 628 abdominal MRI scans (310 T1W and 318 T2W) collected from patients with liver cirrhosis. The dataset includes manual segmentations of cirrhotic livers across multiple stages of disease progression. All images were acquired in clinical settings and have undergone full anonymization. Our data collection adhered to the following protocol: (1) Scanner diversity: MRI scans were obtained from three different scanner models to ensure dataset heterogeneity: Philips Achieva 1.5T, Philips Achieva 3T, and Siemens Symphony 1.5T. (2) Image Selection Criteria: All included images met quality standards suitable for radiological assessment. Images with poor quality or significant motion artifacts were excluded, though the dataset does contain scans with mild to moderate artifacts (motion or susceptibility) to represent real-world clinical conditions. (3) Contrast Enhancement: Approximately 95% of T1W images were acquired during the post-contrast portal venous phase to enhance organ-to-vessel contrast. (4) Patient Population: The dataset includes images from patients diagnosed with liver cirrhosis exhibiting various morphological alterations, including: Contour nodularity, Hepatic segment atrophy or hypertrophy, Complications such as ascites, varices, and splenomegaly. (5) Control Group: A smaller set of non-cirrhotic control subjects (n=55) is included for comparison purposes.
The CirrMRI600+ dataset includes several metadata files to facilitate data usability and provide relevant subject information. The primary file, "…CompleteData-age-gender-evaluation.csv," contains demographic information (age, gender) and radiological (visual) evaluations of cirrhosis (1: Mild, 2: Moderate, 3:Severe) for all 337 patients included in the study. For researchers focusing specifically on T1W or T2W imaging, we provide separate files: "T1-age-gender-evaluation.csv" details the demographic and evaluation data for the 310 patients with "T1-weighted MRI scans", while "T2-age-gender-evaluation.csv" provides the same information for the 318 patients with T2W scans. To support paired-image analysis, the "…-Paired-age-gender-evaluation.csv" file identifies the 291 patients who have both T1W and T2W scans available. Additionally, the dataset includes "Healthy-demographics.csv," which documents age and gender information for the 55 individuals in the control group. To ensure proper interpretation of all variables, we provide a "Labels.txt" file that defines all column headings, coding schemes, and measurement units used throughout the metadata files.
Technical Validation
Validation of Data Collection.
All images underwent quality assessment by participating radiologists prior to inclusion. Manual segmentations were performed by experienced radiologists following a standardized protocol as mentioned earlier in Methods section. The annotations provide ground truth segmentations of cirrhotic livers for all included MRI scans. We also provided radiological evaluation of each patient’s MRI scans for cirrhosis severity (1: Mild, 2: Moderate, 3: Severe). This evaluation was done by our participating radiologist(s) based on organ volume and shape, decompensation and complication status (ascites, splenomegaly, HCC, varices), surface nodularity, and parenchymal heterogeneity-enhancement.
Validation of baseline segmentation models.
We carefully selected several baseline methods from CNN-family (e.g., VNet [26], AttentionU-Net [27]) and from transformer-based methods (for example, nnUNet [35], Swin UNetR [33], nnFormer3D [34], LinTransUNet [32] and TransUNet3D [31]. All models were implemented in PyTorch 2.2.2 with CUDA 11.2 on dual Nvidia A6000 GPUs (48GB each) using PyTorch’s Distributed Data Parallel. For 3D models, we employed a BCE-Dice loss function with AdamW optimizer, initial learning rate of 0.0001 with Cosine Annealing Scheduler, and decay rate of 0.001 every 10 epochs. Each volume was standardized to dimensions and processed with a batch size of 4. The 2D models maintained similar optimization parameters but used a batch size of 16 with the Adam optimizer. All networks were trained for up to 500 epochs with early stopping (patience=50) to prevent overfitting. Model performance was evaluated using multiple metrics: mIoU, Dice coefficient, HD95, precision, recall, and ASDD. To avoid overfitting and provide smooth training, we used rotation, translation, scaling, shear, and intensity transformation (shift, contrast adjustment, noise addition) to enhance the robustness and generalizability of segmentation methods.
@@
Technical Validation of Segmentation Masks on CirrMRI600+T1W: Table 2 presents a comprehensive evaluation of 11 SOTA 3D segmentation networks on the CirrMRI600+T1W dataset. As summarized, nnSynergyNet3D achieved the highest overall performance with mIoU of 84.51, DSC of 87.89%, HD95 of 21.04 mm, and precision of 88.72%. nnSynergyNet3D performed better because of its synergistic and auto-configured continuous and discrete representation, allowing the model to capture fine and coarse features along with long-range dependencies due to its Transformer-inspired design. LinTransUnet and nnFormer3D demonstrated comparable performance in capturing cirrhotic liver tissue and its boundaries. Their performances were attributed to their Transformer based design with auto configuration, which enabled the models to learn and adapt to the liver’s varying shape and complex boundaries. Again, this highlighted the importance of long-range dependencies. Conversely, nnUNet3D demonstrated slightly poorer performance, underscoring the significance of Transformer-based representations for cirrhotic liver segmentation. Models like SwinUNeTr, TransBTS, and TransUNet3D do not significantly surpass CNN-based models such as nnUNet and SynergyVNet3D, showing the importance of auto-configuration and hybrid CNN-Transformer-based models.
Technical Validation of Segmentation Masks on CirrMRI600+T2W: Table 3 presents evaluation of SOTA 3D segmentation networks on the CirrMRI600+T2W dataset. We observed similar results to those in the T1W segmentation results. nnSynergyNet3D has a superior DSC value of 86.51%, the lowest HD of 24.19 mm, and the lowest ASDD value of 3.96 mm. nnFormer3D and nnUNet3D are the other two competitive networks. The models such as SwinUNeTr, TransBTS, and TransUNet3D do not significantly surpass CNN-based models such as nnUNet and SynergyVNet3D, emphasizing the importance of auto-configured and hybrid CNN-Transformer-based models for achieving competitive segmentation results. It should be also note that the T2W images were predominantly acquired using accelerated techniques such as half-fourier single shot turbo-spin-echo (HASTE). The T2W images were acquired in a slice-by-slice fashion in the axial plane (the physics of HASTE generates pseudo-3D volumes). Due to the long echo times required for T2W image acquisition, motion artifacts in the craniocaudal plane from diaphragm movement during respiration are expected. Due to this, we trained T2W-based segmentation models using both volumetric data and 2D planar data. Additionally, we show the comparisons between the different models. By including the experiments on the 2D dataset, we want to develop a robust algorithm that could work well with motion artifacts that might have affected the 3D volumetric scans. We believe that for segmentation models using T2W images, analysis and segmentation using 2D images is an appropriate choice given the expected motion artifacts arising from the acquisition techniques.
@@
Qualitative Validation of Segmentation Masks on both modalities: Fig. 2 and Fig. 3 show qualitative results for T1W and T2W samples, respectively. These results demonstrated that existing models perform well in segmenting cirrhotic liver under mild conditions. As highlighted by white boundaries, these models suffer under moderate-to-severe cases due to the poor texture of MRIs caused by cirrhosis scarring. nnSynergyNet3D consistently performed well even for advanced cirrhotic livers compared to other SOTAs such as nnUNet, TransUNet, and Attention-UNet. This is likely because of its auto-configured, hybrid, and synergistic nature, enabling it to capture the texture of the cirrhotic liver more effectively. Figure 4 shows the qualitative results of different methods on CirrMRI600+T2W datasets. From the qualitative results, it can be observed that models such as UNet produce under-segmentation, whereas nnUNet produces over-segmentation. SynergyNet also produces shows under and over-segmentation for diverse cases. However, MedSegDiff is better at handling complexity. The team of four clinicians verified all the segmented volumetric scans. We obtained high inter-observer agreement (kappa scores of 0.89 for T1W and 0.87 for T2W).
@@
Consistency Analysis of CirrMRI600+ Masks with Benchmarked Methods: We established six baseline models for our CirrMRI600+T2W 2D dataset, spanning diverse architectures: UNet, AttentionUNet, nnUNet2D (with deep supervision), TransUNet (transformer-based), SynergyNet, and MedSegDiff (diffusion-based) (See Table 4). This range covers key approaches in medical image segmentation. Quantitative comparisons revealed SynergyNet as the top UNet-based performer, achieving a mIoU of 0.7383, Dice coefficient of 0.7592, HD95 of 30.94, precision of 0.7882, recall of 0.8222, and ASDD of 7.55. However, MedSegDiff outperformed all methods with a mIoU of 0.7489, Dice coefficient of 0.7667, HD95 of 30.89, and ASDD of 7.34, while SynergyNet maintained higher precision and recall values.
Technical Challenge-Single Center Study: While CirrMRI600+ offers a significant leap forward in liver cirrhosis segmentation research, it’s important to acknowledge its limitations. One such limitation is its single-center nature. Ideally, a truly comprehensive dataset would encompass data from multiple medical institutions, capturing the potential variations in acquisition protocols and patient populations encountered in real-world clinical settings. However, CirrMRI600+ addresses this limitation by prioritizing other key strengths. The dataset was very carefully curated, ensuring the highest quality annotations and ground truth segmentation for all included scans. Furthermore, CirrMRI600+ boasts a rich heterogeneity in disease states within the cirrhotic population. By carefully selecting a diverse range of cirrhotic liver presentations, the dataset ensures generalizability to a broader spectrum of cirrhosis cases.
Technical Challenge-Not all T1Ws are contrast-enhanced: Another weakness was that more than 95% of the T1W images were contrast enhanced and not all possible contrast types and phases of contrast administration were represented in the dataset. Although the cross-modality testing shows strong generalizability, small annotations might have been missed, and our annotations might not be correct pixel precise for all the cases. Other modalities, such as computed tomography (CT) scan, ultrasound, elastography, and magnetic resonance elastography, were absent besides MRI.
@@
Vision for Multi-Organ Annotations: To ensure high-quality annotations while maintaining efficiency, CirrMRI600+ leverages a semi-automated approach in ground truth construction. Deep learning model predictions served as a starting point, followed by refinement from radiologists. This approach enabled radiologists to focus on potential algorithm failures and refine the output masks more efficiently, saving valuable annotation time and cost. Rigorous benchmarking using the nnSynergyNet3D method on both T1W and T2W scans demonstrates promising performance, achieving Dice Similarity Coefficients (DSC) of 87.89% and 86.51%, respectively. CirrMRI600+ lays the foundation for a future multi-organ dataset. Our long-term vision is to expand the annotations to include additional organs such as kidneys, spleen, pancreas, and major vessels, creating a comprehensive resource for abdominal organ segmentation tasks for multimodal MRIs.
Other Technical Challenges and Potential Solutions: While our benchmark results demonstrate strong performance for liver segmentation (achieving DSC of 87.89% for T1W and 86.51% for T2W with nnSynergyNet3D), significant challenges and research opportunities remain in the domain of cirrhotic liver analysis. We identify several key areas for future investigation using the CirrMRI600+ dataset:
Segmentation: Despite the promising results, our qualitative analysis reveals persistent difficulties in accurately delineating certain anatomical regions, particularly in cases with severe cirrhosis. These challenges include:
(1) Boundary ambiguity in advanced cirrhosis: The irregular, nodular boundaries characteristic of advanced cirrhotic livers remain difficult to precisely segment, especially where the liver interfaces with adjacent structures having similar intensity profiles.
(2) Segmentation of heterogeneous parenchyma: The variable signal intensities within severely fibrotic regions often lead to under-segmentation of areas that deviate significantly from typical liver appearance.
(3) Handling of focal lesions: The presence of focal lesions (including regenerative nodules and potential hepatocellular carcinoma) introduces additional complexity that current methods struggle to consistently address.
(4) Inter-modality performance gap: The observed differences in segmentation performance between T1W and T2W sequences suggest opportunities for developing specialized approaches that better leverage the unique characteristics of each modality.
Usage Notes
The entire dataset can be downloaded from the OSF repository [25]. To process the provided images and segmentation maps (for all the compared benchmarking methods), we highly recommended using medical imaging tools such as ITKSnap, 3D-Slicer, Amira, or similar dicom viewers. We verified that all NIfTI files (segmentations and images) and DICOM files can be loaded correctly with 3D-Slicer (https://www.Slicer.org) and ITKSnap.
Code Availability
All the segmentation codes, model files, and customized instructions are available to be downloaded from our dedicated GitHub repository (https://github.com/NUBagciLab/CirrMRI600Plus) [36].
Acknowledgements
This project is supported by NIH funding: R01-CA246704, R01-CA240639, U01-DK127384-02S1, and U01-CA268808.
Author contributions statement
A. Medetalibeyoglu, G. Durak, D. Jha and U. Bagci conceptualized the work. U. Bagci, G. Durak and A. Medetalibeyoglu were critical in designing the study and facilitating the data collection. D. Jha, O. K. Susladkar and V. Gorade performed the baseline experiments. G. Durak led the data annotation process coordinating effort with a team including E. Keles, M. Antalek, D. Seyithanoglu, T. Cebeci, H. Ertugrul Aktas, G. D. Kartal, S. Kaymakoglu, S. M. Erturk, Y. Velichko, D. Ladner, A. A. Borhani, and A. Medetalibeyoglu. All the images were reviewed by the radiology team first, and consensus was achieved with the clinical team about the annotation style, time, repetition, and other clinical questions related to cirrhosis severity estimation. T. Cebeci, G. Kartal, S. Kaymakoglu, and S. Erturk helped in data collection, IRB, and secure transfer of the files. All authors contributed to the paper’s writing and revision.
Code statement
The code for baseline methods are available at https://github.com/NUBagciLab/CirrMRI600Plus.
Competing interests
The authors declare no competing interests regarding this publication and the dataset.